Shift report 20020108 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 08 2002 - 08:22:46 EST

  • Next message: Flemming Videbaek: "Fw: Electron Ion Collider Registration"

    ------------------------------------------------------------
    Date: Tue Jan  8 08:22:46 2002, Shift: 00:00-08:00
    Supervisor: ykl, with: 
    Log book#12 pages 23 - 26
    
    Shift Summary:
    ------------------------------------------------------------
    1. Relatively low intensity beam for half of the time.
    
    2. Runs 6344 (10 minutes) and 6345 (15 minutes) are for
    testing the INEL trigger. The first rings (1L and 1R) were
    removed from the trigger for this test.
    
       Run 6346 has Trigger 3(MRS) 95K    @40 deg 4B
                    Trigger 6(FS) 46K     @12 deg 1/8B
    
       Run 6353 has been running for an hour and half with
                   the same spectrometer settings and is
                   still on.
    
    3. Tripping of D1U2 (which happened several times) is
    difficult to monitor. Tripping was missed twice.
    
    D3 tripped once.
    
    T4 anode tripped once.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 08 2002 - 08:23:26 EST